A 1500 W electric geyser used every day for 2hrs . Calculate the energy consumed ?
 

Grade:12th pass

1 Answers

Gaurav
3 years ago
Given,
Power of geyser is 1500
& Hours of use is 2 hrs=60×60=3600s
So,
Total consume energy is=1500×2×3600
                                           =10800000 joule